OpenStudy (anonymous):

how do i solve this? Write an equation of an ellipse centered at the origin with the height 8 units and width 16 units.

OpenStudy (mathmale):

the standard eq'n of an ellipse is\[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] are y ou able to write such an equation for the sit'n where the center is at (h,k)?

OpenStudy (mathmale):

You don't absolutely have to answer that, since the given ellipse is centered at the origin. What do you suppose "height" means in this context? Is it the full minor axis of the ellipse, or just half of the minor axis?

OpenStudy (mathmale):



OpenStudy (anonymous):

i am so confused